# Robustness in Large Language Models: A Survey of Mitigation Strategies and Evaluation Metrics

### v1

@@ 3.2 Training Related Sources
Standard pre-training objectives like next-token prediction and fine-tuning objectives based on Empirical Risk Minimization (ERM) primarily optimize for average performance on the training distribution. These objectives may not explicitly encourage robustness to distribution shifts or perturbations. Furthermore, the dynamics of gradient-based optimization might lead models to latch onto simple, non-robust features early in training. The process of training and aligning LLMs can introduces such vulnerabilities:
3.2.1 Optimization Objectives (ERM)
The standard training objective, Empirical Risk Minimization (ERM), aims to minimize average loss on the training data. On large, biased datasets, instead of spending capacity in learning more complicated, generalizable features, ERM naturally encourages models to locate the simplest prediction patterns, which often are the least reliable shortcuts (Nagarajan et al., 2024).
@@ 3.2.2 RLHF
While Reinforcement Learning from Human Feedback (RLHF) is crucial for making LLMs safer and more helpful, the process itself can introduce robustness Sources in the following ways:
  • Reward Hacking: LLMs may learn to exploit weaknesses or biases in the reward model (RM) used during RLHF, maximizing the reward signal without genuinely fulfilling the intended preference (e.g., generating overly verbose responses because the RM implicitly favours length) (Yan et al., 2024b).
  • Bias Amplification: RLHF optimizes based on human feedback, but this feedback can itself be biased, inconsistent, or fail to capture the full spectrum of desirable behaviour. Optimizing for average preferences might even strengthen covert biases not explicitly penalized by the feedback (Barnhart et al., 2025).
  • Capability Trade-offs: Safety alignment, often achieved via RLHF or related techniques, has been observed to sometimes degrade specific capabilities like complex reasoning (Huang et al., 2025).
@@ 3.2.3 Alignment Tax
This phenomenon, termed the “alignment tax”, means that improving alignment (e.g., helpfulness, harmlessness) might come at the cost of reduced performance on general knowledge benchmarks or can sometimes negatively impact model calibration (leading to overconfidence) or introduce new vulnerabilities (Chen et al., 2020; Yang et al., 2024b). Fine-tuning LLMs for specific tasks or aligning them with human preferences (e.g., using RLHF) can lead to a degradation of capabilities learned during pre-training (e.g., Masked Language Modeling vs. Standard auto-regressive prediction), which can influence downstream behaviour (Lin et al., 2024b). This creates a difficult trade-off for developers.
@@ 3.2.4 Fine-tuning Limitations
While fine-tuning adapts LLMs to specific tasks or domains, it can also reduce robustness, as in cases like:
  • Overfitting: Models can overfit to the specific style, domain, or spurious correlations within the fine-tuning data, harming OOD generalization (Yuan et al., 2023).
  • Noise Sensitivity: Fine-tuning performance is highly sensitive to noise in the instruction-following or preference data (Luo et al., 2024). Even moderate noise levels can cause significant performance drops.
  • Instruction Brittleness: Instruction-tuned models can show substantial performance degradation when test-time instructions are phrased differently from those seen during fine-tuning, indicating a lack of robustness to instruction variations (Aljohani et al., 2025). Specialized models might be even more fragile in this regard.
@@ 3.3 Architectural Limitations
While less explored in the provided sources, the inherent properties of the dominant Transformer architecture, such as the attention mechanism’s focus or the lack of explicit symbolic reasoning modules, might contribute to certain robustness vulnerabilities. The lack of built-in mechanisms for reliable uncertainty quantification or causal reasoning could also be a factor (Zhang, 2023). The underlying architecture of LLMs, predominantly the Transformer, also presents inherent vulnerabilities:
  • Transformer Properties: Specific components of the Transformer architecture, such as the self-attention mechanism, can be points of failure. While properties like redundancy might offer some baseline resilience compared to older architectures, this resilience is limited and can be overcome by targeted attacks Das et al. (2025). Further, expanding model depth (layer expansion) can particularly disrupt the attention mechanisms.
  • Vulnerabilities to Modifications: Practical deployment often requires model modifications for efficiency, such as quantization (reducing numerical precision) or pruning (removing weights/neurons). These compression techniques can induce “information loss cascades”, significantly degrading robustness (Ma et al., 2025). Similarly, architectural changes aimed at efficiency, like attention-efficient variants (e.g., GLA Transformer, MatMul-Free LM (Fan & Tao, 2024) or Funnel Transformers (Choi et al., 2025)), involve trade-offs that may impact robustness. Transformer architectures appear to have inherent robustness thresholds beyond which modifications cause severe performance degradation. For example, the extreme vulnerability to Bit-Flip Attacks (BFAs), where flipping just a few bits in the weights of a large model can cause catastrophic failure, challenges assumptions about their inherent resilience (Ma et al., 2025).
  • Linear Transformers: Simpler variants like linear transformers have shown specific vulnerabilities, such as susceptibility to hijacking attacks during in-context learning, where malicious examples in the prompt manipulate the learned function (Anwar et al., 2024).
@@ 3.4.1 Inference and Decoding Vulnerabilities
  • Decoding Strategy Sensitivity: The algorithm used to select the next token based on the model’s probability distribution (e.g., greedy search, beam search, top-k/nucleus sampling) has a major impact on the performance of the model (Nik et al., 2025).
    • Performance vs Robustness: Different strategies offer trade-offs. Deterministic methods like greedy search (Bang et al., 2023) might be preferred for closed-ended tasks requiring precision, while stochastic methods (sampling) are often better for open-ended generation requiring diversity (Shi et al., 2024a). However, stochasticity can reduce the consistency of the model.
    • Hyperparameter Sensitivity: The performance and behaviour of decoding strategies are often highly sensitive to hyperparameters like temperature (for sampling) or beam width (for beam search). Achieving optimal performance might require extensive tuning, but fixed hyperparameters might lead to suboptimal robustness across diverse inputs (Shi et al., 2024a).
  • Calibration Errors: LLM confidence scores (often derived from output probabilities) frequently fail to accurately reflect the true likelihood of the generated output being correct (Yao et al., 2024). Models, especially those fine-tuned with RLHF, tend to be overconfident. This miscalibration hinders reliable decision-making based on model outputs.
@@ 3.4.2 Retrieval-Augmented Generation (RAG) Sources
RAG systems enhance LLMs by retrieving external information, but their robustness depends heavily on the quality and relevance of the retrieved context (Shen et al., 2024). If the retriever returns inaccurate, irrelevant, or noisy information (due to its own lack of robustness), the final LLM output quality can be significantly compromised, sometimes performing worse than without retrieval at all. LLMs may also struggle to effectively utilize a large number of retrieved documents, even with long context windows (Yu et al., 2024).
@@ 4 Mitigation Strategies
The strategies are categorized based on the stage of the typical LLM development and deployment pipeline where they are applied:
@@
This pipeline-based categorization provides a structured way to understand where different interventions fit within the LLM lifecycle and how they contribute to overall robustness. Figure 4 provides an illustration of enhancing the robustness of LLMs using these strategies throughout the pipeline.
@@ 4.1 Pre-processing strategies
Data augmentation aims to increase the diversity and size of the training dataset, typically to improve model generalization. Dong et al. (2021) describes this as specifically designed techniques to expose the model to the types of variations or adversarial inputs it might encounter during deployment. Early techniques adapted from general NLP, such as Easy Data Augmentation (EDA), involved simple lexical operations like synonym replacement, random insertion, random swap, or random deletion (Weng, 2023). While useful for generalization, these methods may not adequately prepare models for targeted adversarial attacks or significant distribution shifts.
@@
A more targeted approach is Adversarial Data Augmentation (ADA), which generates challenging examples to improve robustness against data shifts or corruptions. ADA creates misleading target distributions using adversarial loss, perturbing data to fool the model during training. A key advancement, Maximum-Entropy ADA (ME-ADA) (Zhao et al., 2020), introduced an Information Bottleneck-derived regularizer that maximizes model uncertainty, producing harder adversarial examples by pushing augmented data further from the source distribution, outperforming prior methods on benchmarks. Simultaneously, Adversarial Contrastive Learning (ACL) (Jiang et al., 2020) merges adversarial examples with self-supervised pre-training. By enforcing feature consistency across standard and adversarial views, ACL enhances inherent invariance, improving robustness and label efficiency over unsupervised methods. Follow-up work refined ACL via robustness-aware coreset selection and adversarial invariant regularization (Xu et al., 2023).
@@
As research shifted towards LLMs, augmentation techniques became more specialised. Liu & Sun (2023) introduced the Adversarial Augmentation Approach (A3), combining adversarial training with NLP-specific data augmentation. Unlike traditional methods, A3 employs a paraphrasing model guided by a separate generator to produce reusable adversarial examples, optimized for specific tasks via a discriminator. This efficiency boosted accuracy for models like BERT, outperforming earlier methods. Further, Bae et al. (2025) introduces SALAD (Structure Aware and LLM- driven Augmented Data), a contrastive learning approach that employs a tagging-based method for generating structure-aware positive samples and leverages LLMs to create diverse counterfactual negative samples for triplet loss optimization.
@@
Meanwhile, HarmAug (Lee et al., 2024b) tackles safety: it trains compact “guard” models to detect harmful queries by generating adversarial examples (e.g., jailbroken LLM prompts) and distilling knowledge from a large teacher model. This let a 435M-parameter safety model match the performance of 7B+ counterparts, proving targeted augmentation can compress robustness into smaller systems. These advances align with broader efforts to refine LLM robustness through data-centric strategies, such as adaptive data filtering and synthetic data generation (Dong et al., 2021; Qian et al., 2022; Ding et al., 2024; Zayed et al., 2022; Lim et al., 2023).
@@
Modern large language models (LLMs) rely heavily on vast amounts of internet-sourced data for training, but this data often includes noise, toxic content, societal biases, personal information, and factual inaccuracies (Huang et al., 2024a). Since LLMS can memorise such flaws during training, these issues directly translate into risks like biased outputs, privacy breaches, and unreliable behaviour in real-world applications. Data filtering approaches this by modifying training samples or their weights to mitigate harmful content in model learning (Li et al., 2025b; Wang et al., 2025). Additionally, this approach can generate contrastive pairs (e.g., pairing biased and debiased examples) to explicitly teach models to disregard biased patterns in the data. While filtering and cleaning data before training is essential, considering the fact that manual curation is time-consuming and demands domain expertise, to mitigate these risks, it also introduces complex trade-offs: overly strict filtering to remove toxic content or reduce bias can inadvertently harm the model’s general performance or erase valid examples from underrepresented groups. For instance, prioritizing safety might suppress harmful outputs but weaken the model’s versatility, while aggressive bias mitigation could strip away nuanced data critical for fairness. Thus, achieving the right balance between data quality (removing harmful content) and quantity (retaining useful diversity) remains a pivotal challenge in building trustworthy LLMs.
@@
Laurençon et al. (2023) proposed a Rule-based Heuristic Filtering, using predefined criteria to remove sensitive data (e.g., PII) or toxic content via blocklists or classifiers, though their success hinges on the precision of these tools. To streamline complexity, frameworks like Data-Juicer (Chen et al., 2023a) systematize cleaning by automating steps like deduplication and quality scoring. More recently, LLMs themselves have been repurposed as data cleaners: studies show they can fix simple errors (e.g., invalid dates) using contextual analysis but falter with nuanced issues like dataset-wide biases (Bendinelli et al., 2025). From basic heuristics to AI-driven refinement, these approaches illustrate the layered effort to balance scalability and reliability in training data. Further, BaichuanSEED (Dong et al., 2024a) demonstrates the scalability of data-centric methods, showing that careful data curation and reweighting can enable even a general 7B-parameter LLM to achieve state-of-the-art benchmark performance without architectural modifications. Yang et al. (2024c) introduces RAZOR (Rewriting And Zero-bias Optimization Refinement), a novel unsupervised text rewriting technique designed to mitigate dataset biases (shortcuts) that hinder language model generalization without requiring prior knowledge of specific biases. RAZOR iteratively rewrites text segments using LLMs, selecting replacements that reduce spurious correlations between surface features and labels based on token statistics and positional information.

### v2

@@ 3.2 Training Related Sources
Standard pre-training objectives like next-token prediction and fine-tuning objectives based on Empirical Risk Minimization (ERM) primarily optimize for average performance on the training distribution. These objectives may not explicitly encourage robustness to distribution shifts or perturbations. Furthermore, the dynamics of gradient-based optimization might lead models to latch onto simple, non-robust features early in training. The process of training and aligning LLMs, therefore, introduces several potential vulnerabilities, including:
3.2.1 Optimization Objectives
LLMs are primarily trained using next-token prediction loss (NTP), which optimizes the likelihood of predicting each token in a sequence given its preceding context. This objective drives the model to capture patterns that reduce sequence-level prediction loss, enabling the acquisition of syntactic, semantic, and contextual dependencies. But it can also encourage reliance on superficial patterns or shortcuts in the training data, potentially limiting generalization and robustness (thrampoulidis2024implicit; nagarajan2024understandingfailuremodesoutofdistribution).
@@ 3.2.2 RLHF
While Reinforcement Learning from Human Feedback (RLHF) is crucial for making LLMs safer and more helpful, the process itself can hinder robustness in the following ways:
  • Reward Hacking: LLMs may learn to exploit weaknesses or biases in the reward model (RM) used during RLHF, maximizing the reward signal without genuinely fulfilling the intended preference (e.g., generating overly verbose responses because the RM implicitly favours length) (yan2024rewardrobustrlhfllms).
  • Bias Amplification: RLHF optimizes based on human feedback, but this feedback can itself be biased, inconsistent, or fail to capture the full spectrum of desirable behaviour. Optimizing for average preferences might even strengthen covert biases not explicitly penalized by the feedback (barnhart2025aligningwhatlimitsrlhf).
  • Capability Trade-offs: Safety alignment, often achieved via RLHF or related techniques, has been observed to sometimes degrade specific capabilities like complex reasoning (huang2025safetytaxsafetyalignment).
@@ 3.2.3 Alignment Tax
This phenomenon, termed the “alignment tax”, means that improving alignment (e.g., helpfulness, harmlessness) might come at the cost of reduced performance on general knowledge benchmarks or can sometimes negatively impact model calibration (leading to overconfidence) or introduce new vulnerabilities (chen2020distillingknowledgelearnedbert; yang2024confidencecalibrationrationalizationllms). Fine-tuning LLMs for specific tasks or aligning them with human preferences (e.g., using RLHF) can lead to a degradation of capabilities learned during pre-training (e.g., Masked Language Modeling vs. Standard auto-regressive prediction), which can influence downstream behaviour (lin2024mitigatingalignmenttaxrlhf). This creates a difficult trade-off for developers.
@@ 3.2.4 Fine-tuning Limitations
While vanilla fine-tuning adapts LLMs to specific tasks or domains, it can also reduce robustness (sengupta2025robustefficientfinetuningllms; luo2024robustftrobustsupervisedfinetuning), as in cases like:
  • Overfitting: Models can overfit to the specific style, domain, or spurious correlations within the fine-tuning data, harming OOD generalization (yuan2023revisiting).
  • Noise Sensitivity: Fine-tuning performance is highly sensitive to noise in the instruction-following or preference data (luo2024robustftrobustsupervisedfinetuning). Even moderate noise levels can cause significant performance drops.
  • Instruction Brittleness: Instruction-tuned models can show substantial performance degradation when test-time instructions are phrased differently from those seen during fine-tuning, indicating a lack of robustness to instruction variations (aljohani2025comprehensivesurveytrustworthinesslarge). Specialized models might be even more fragile in this regard.
@@ 3.3 Architectural Limitations
While less explored in the provided sources, the inherent properties of the dominant Transformer architecture, such as the attention mechanism’s focus or the lack of explicit symbolic reasoning modules, might contribute to certain robustness vulnerabilities. The lack of built-in mechanisms for reliable uncertainty quantification or causal reasoning could also be a factor (zhang2023awesome). The underlying architecture of LLMs, predominantly the Transformer, also presents inherent vulnerabilities:
  • Transformer Properties: Specific components of the Transformer architecture, such as the self-attention mechanism, can be points of failure. While properties like redundancy might offer some baseline resilience compared to older architectures, this resilience is limited and can be overcome by targeted attacks das2025genbfaevolutionaryoptimizationapproach. Further, expanding model depth (layer expansion) can particularly disrupt the attention mechanisms.
  • Vulnerabilities to Modifications: Practical deployment often requires model modifications for efficiency, such as quantization (reducing numerical precision) or pruning (removing weights/neurons). These compression techniques can induce “information loss cascades”, significantly degrading robustness (ma2025modelhemorrhagerobustnesslimits). Similarly, architectural changes aimed at efficiency, like attention-efficient variants (e.g., GLA Transformer, MatMul-Free LM (fan2024resilientefficientllmscomparative) or Funnel Transformers (choi2025revisitingfunneltransformersmodern)), involve trade-offs that may impact robustness. Transformer architectures appear to have inherent robustness thresholds beyond which modifications cause severe performance degradation. For example, the extreme vulnerability to Bit-Flip Attacks (BFAs), where flipping just a few bits in the weights of a large model can cause catastrophic failure, challenges assumptions about their inherent resilience (ma2025modelhemorrhagerobustnesslimits).
  • Linear Transformers: Simpler variants like linear transformers have shown specific vulnerabilities, such as susceptibility to hijacking attacks during in-context learning, where malicious examples in the prompt manipulate the learned function (anwar2024adversarialrobustnessincontextlearning).
Figure 4: An illustration of some failure cases of non-robustness (adapted from schulhoff2025promptreportsystematicsurvey; shen-etal-2024-assessing; gallegos2024biasfairnesslargelanguage; bano2025doessoftwareengineerlook; zhou-etal-2024-explore). The examples are representative of behaviors primarily observed in GPT-3.5, and are intended to demonstrate common categories of robustness failures.
@@ 3.4.1 Inference and Decoding Vulnerabilities
  • Decoding Strategy Sensitivity: The algorithm used to select the next token based on the model’s probability distribution (e.g., greedy search, beam search, top-k/nucleus sampling) has a major impact on the performance of the model (nik2025energyconsciousllmdecodingimpact).
    • Performance vs Robustness: Different strategies offer trade-offs. Deterministic methods like greedy search (bang2023multitaskmultilingualmultimodalevaluation) might be preferred for closed-ended tasks requiring precision, while stochastic methods (sampling) are often better for open-ended generation requiring diversity (shi2024thoroughexaminationdecodingmethods). However, stochasticity can reduce the consistency of the model.
    • Hyperparameter Sensitivity: The performance and behaviour of decoding strategies are often highly sensitive to hyperparameters like temperature (for sampling) or beam width (for beam search). Achieving optimal performance might require extensive tuning, but fixed hyperparameters might lead to suboptimal robustness across diverse inputs (shi2024thoroughexaminationdecodingmethods).
  • Calibration Errors: LLM confidence scores (often derived from output probabilities) frequently fail to accurately reflect the true likelihood of the generated output being correct (yao2024retroformerretrospectivelargelanguage). Models, especially those fine-tuned with RLHF, tend to be overconfident. This miscalibration hinders reliable decision-making based on model outputs.
@@ 3.4.2 Retrieval-Augmented Generation (RAG) Sources
RAG systems enhance LLMs by retrieving external information, but their robustness depends heavily on the quality and relevance of the retrieved context (shen-etal-2024-assessing). If the retriever returns inaccurate, irrelevant, or noisy information (due to its own lack of robustness), the final LLM output quality can be significantly compromised, sometimes performing worse than without retrieval at all. LLMs may also struggle to effectively utilize a large number of retrieved documents, even with long context windows (yu2024rankragunifyingcontextranking).
@@ 4 Mitigation Strategies
Having covered the various sources of non-robustness in Section 3, we will now look into mitigation strategies. These strategies are categorized based on the stage of the typical LLM development and deployment pipeline where they are applied:
@@
This pipeline-based categorization provides a structured way to understand where different interventions fit within the LLM lifecycle and how they contribute to overall robustness. Further, Table 3 summarizes representative methods under each category, while Figure 5 illustrates how these strategies collectively enhance robustness across the LLM development pipeline.
@@ 4.1 Pre-processing strategies
Data augmentation aims to increase the diversity and size of the training dataset, typically to improve model generalization. dong-etal-2021-data describes this as specifically designed techniques to expose the model to the types of variations or adversarial inputs it might encounter during deployment. Early techniques adapted from general NLP, such as Easy Data Augmentation (EDA), involved simple lexical operations like synonym replacement, random insertion, random swap, or random deletion (weng2023attack). While useful for generalization, these methods may not adequately prepare models for targeted adversarial attacks or significant distribution shifts.
@@
A more targeted approach is Adversarial Data Augmentation (ADA), which generates challenging examples to improve robustness against data shifts or corruptions. ADA creates misleading target distributions using adversarial loss, perturbing data to fool the model during training. A key advancement, Maximum-Entropy ADA (ME-ADA) (Zhao_Liu_Peng_Metaxas_2020), introduced an Information Bottleneck-derived regularizer that maximizes model uncertainty, producing harder adversarial examples by pushing augmented data further from the source distribution, outperforming prior methods on benchmarks. Simultaneously, Adversarial Contrastive Learning (ACL) (jiang2020robustpretrainingadversarialcontrastive) merges adversarial examples with self-supervised pre-training. By enforcing feature consistency across standard and adversarial views, ACL enhances inherent invariance, improving robustness and label efficiency over unsupervised methods. Follow-up work refined ACL via robustness-aware coreset selection and adversarial invariant regularization (xu2023efficientadversarialcontrastivelearning). Furthermore, safety propose a data augmentation approach using “safety recovery examples" to cultivate “deep safety alignment", which significantly improves robustness against common exploits like prefilling and adversarial suffix attacks.
@@
As research shifted towards LLMs, augmentation techniques became more specialised. liu-sun-2023-end introduced the Adversarial Augmentation Approach (A3), combining adversarial training with NLP-specific data augmentation. Unlike traditional methods, A3 employs a paraphrasing model guided by a separate generator to produce reusable adversarial examples, optimized for specific tasks via a discriminator. This efficiency boosted accuracy for models like BERT, outperforming earlier methods. Further, bae2025saladimprovingrobustnessgeneralization introduces SALAD (Structure Aware and LLM- driven Augmented Data), a contrastive learning approach that employs a tagging-based method for generating structure-aware positive samples and leverages LLMs to create diverse counterfactual negative samples for triplet loss optimization.
@@
Meanwhile, HarmAug (Lee2024HarmAugED) tackles safety: it trains compact “guard” models to detect harmful queries by generating adversarial examples (e.g., jailbroken LLM prompts) and distilling knowledge from a large teacher model. This let a 435M-parameter safety model match the performance of 7B+ counterparts, proving targeted augmentation can compress robustness into smaller systems. These advances align with broader efforts to refine LLM robustness through data-centric strategies, such as adaptive data filtering and synthetic data generation (dong-etal-2021-data; qian2022perturbationaugmentationfairernlp; ding-etal-2024-data; zayed2022deeplearninghealthydata; 10203320).
@@
Modern LLMs rely heavily on vast amounts of internet-sourced data for training, but this data often includes noise, toxic content, societal biases, personal information, and factual inaccuracies (huang2024trustllmtrustworthinesslargelanguage). Since LLMs can memorise such flaws during training, these issues directly translate into risks like biased outputs, privacy breaches, and unreliable behaviour in real-world applications. Data filtering approaches this by modifying training samples or their weights to mitigate harmful content in model learning (li2025trustworthymachinelearningmemorization; wang2025surveyresponsiblellmsinherent). Additionally, this approach can generate contrastive pairs (e.g., pairing biased and debiased examples) to explicitly teach models to disregard biased patterns in the data. While filtering and cleaning data before training is essential, considering the fact that manual curation is time-consuming and demands domain expertise, to mitigate these risks, it also introduces complex trade-offs: overly strict filtering to remove toxic content or reduce bias can inadvertently harm the model’s general performance or erase valid examples from underrepresented groups. For instance, prioritizing safety might suppress harmful outputs but weaken the model’s versatility, while aggressive bias mitigation could strip away nuanced data critical for fairness. Thus, achieving the right balance between data quality (removing harmful content) and quantity (retaining useful diversity) remains a pivotal challenge in building trustworthy LLMs.
Figure 5: Comprehensive LLM Robustness Pipeline. A multi-stage framework for enhancing LLM robustness across the deployment lifecycle. (1) Pre-processing: Data curation and augmentation strategies; (2) In-processing: Training-time interventions including adversarial training and alignment; (3) Intra-processing: Real-time inference adaptations; (4) Post-processing: Output validation and filtering. Arrows indicate data flow and feedback mechanisms between stages.
@@
laurençon2023bigsciencerootscorpus16tb proposed a Rule-based Heuristic Filtering, using predefined criteria to remove sensitive data (e.g., PII) or toxic content via blocklists or classifiers, though their success hinges on the precision of these tools. To streamline complexity, frameworks like Data-Juicer (chen2023datajuiceronestopdataprocessing) systematize cleaning by automating steps like deduplication and quality scoring. More recently, LLMs themselves have been repurposed as data cleaners: studies show they can fix simple errors (e.g., invalid dates) using contextual analysis but falter with nuanced issues like dataset-wide biases (bendinelli2025exploringllmagentscleaning). From basic heuristics to AI-driven refinement, these approaches illustrate the layered effort to balance scalability and reliability in training data. Further, BaichuanSEED (dong2024baichuanseedsharingpotentialextensive) demonstrates the scalability of data-centric methods, showing that careful data curation and reweighting can enable even a general 7B-parameter LLM to achieve state-of-the-art benchmark performance without architectural modifications. yang2024razorsharpeningknowledgecutting introduces RAZOR (Rewriting And Zero-bias Optimization Refinement), a novel unsupervised text rewriting technique designed to mitigate dataset biases (shortcuts) that hinder language model generalization without requiring prior knowledge of specific biases. RAZOR iteratively rewrites text segments using LLMs, selecting replacements that reduce spurious correlations between surface features and labels based on token statistics and positional information.